OpenStudy (anonymous):

4 times the measure of the complement of an angle is equal to / the measure of the original angle's supplement. What is the measure of the angle? 18° 48° 72° 108°

OpenStudy (anonymous):

is equal to 2/3

OpenStudy (midhun.madhu1987):

Let the angle be x. Complement of x = 90 - x Supplement of x = 180 - x 4(90 - x) = 180 - x Find x from this...

OpenStudy (anonymous):

72%